OpenStudy (anonymous):

Solve: −3x − 3 = 12 5 −5 −3 3

OpenStudy (anonymous):

-3x = 15 x = -5

OpenStudy (anonymous):

where did you get the 15?

OpenStudy (anonymous):

−3x − 3 = 12 -3 (x+1) = 12 x+1 = -4 x=-5

OpenStudy (anonymous):

why did you put the x+1 instead of x-3? am i supposed to change it to that in every algabra expression??

OpenStudy (anonymous):

ok heres a more detailed solution